# Goal-Conditioned Data Augmentation for Offline Reinforcement Learning

### v1

@@ IV-B Selective goal conditions
2:Split initial offline dataset $\mathcal{D}$ into trajectories according to episode terminal information
3:Calculate RTGs for each transition sample in trajectories
4:Add RTGs and timesteps as goals into $\mathcal{D}$
5:Train $G_{\theta}$ on $\mathcal{D}$ using Eq.  9 by conditioning on goals
6:repeat
7:     Extract goal conditions for sampling according to the goal selection mechanism.
@@
8:     Re-assign sampling goal conditions with goal scaling factor $\lambda$ using Eq. 6
@@
9:     Sampling a batch of new transition samples $\mathcal{B}^{*}$
10:     $\mathcal{D}^{*}\leftarrow\mathcal{D}^{*}\cup\mathcal{B}^{*}$
11:until end
12:Train policy $\pi$ on the final dataset $\mathcal{D}^{*}\cup\mathcal{D}$
IV-C Controllable goal scaling
Selective goal conditions offer high-return guidance during the sampling process but are limited in generating data with returns or quality beyond the initial offline datasets. To overcome this limitation, we introduce a controllable goal scaling factor, $\lambda$, which can be multiplied with the goal values to represent a higher return expectation. This approach enables flexible adjustment of goal values to drive the sampling process toward higher-quality data. As illustrated in Figure 1, a higher RTG goal at each timestep directs the sampling process toward a trajectory with a greater overall return. Figure 3 depicts that higher goals guide the sampling process towards higher-reward data regions. Since RTG values can be either positive or negative in certain tasks, we propose multiplying positive goals by the scaling factor and dividing negative goals by it.
It is worth noting that a goal increment factor is not applicable to RTG-based goals, since incrementing all RTGs for a trajectory simply means only incrementing the reward at first timestep while keeping all remaining rewards unchanged.
IV-D Adaptive Gated Conditioning
IV-D1 Condition Embedding
The noise level $\sigma$ for diffusion is encoded using Random Fourier Feature [30] embedding. The RTG is processed with a linear transformation to get a hidden representation. The timestep of each RTG is embedded with Sinusoidal positional embedding [31]. We concatenate the RTG and timestep embeddings to form the goal condition, which is then element-wise added to the noise level embedding and used as the conditional input.
IV-D2 Adaptive Gated Long Skip Connection
As shown in the left part of Figure 2, we adopt a long skip connection similar to U-Net [32] to connect MLP blocks at different levels. To capture different information with varying importance weights, we propose a novel adaptive gated long skip connection structure by adding the previous information with an adaptive gate mechanism based on the given conditions.
$x_{\text{out}}=(1-\omega)*x_{\text{skip}}+\omega*x,$ (7)
where $x_{\text{skip}}$ and $x$ are outputs of a shallower and the previous block, and $\omega$ denotes a learnable weight calculated by regressing the conditional input with an MLP and a sigmoid layer.
IV-D3 Gated Residual MLP Block
The right part of Figure 2 depicts the structure of each gated residual MLP block. We adopt the widely used adaptive layer normalization (adaLN) method [33] to learn dimension-wise scale $\gamma$ and shift $\beta$ based on the conditional information. Besides, we explore a modification of the residual connection [34] and introduce a novel condition-adaptive gated residual connection to further enhances the model by selectively preserving input information. It also regresses the conditional input and gets a learnable weight $\nu$ for adaptively preserving input information.
$x_{\text{out}}=(1-\nu)*F(x)+\nu*x,$ (8)
where $F$ is the learned transformation.
IV-E Model Implementation
@@
Given the strong ability of diffusion models to capture complex data distribution and generate high-dimension data, we adopt EDM [15] as our generative model for augmenting offline data. The neural network equipped with adaptive gated conditioning as illustrated in Figure 2 is used as the denoiser function. We train the generative model to approximate the data distribution of the offline dataset and use every transition tuple as a training sample. Given the non-sequence input format, we do not consider complicated structures, e.g., attention mechanisms, but use simple MLPs to process inputs. Algorithm 1 shows the learning process of our GODA method. With goal conditions $\mathbf{c}$ and transition data $\mathbf{x}$ from the original datasets, the generative model $G_{\theta}$ with a learnable denoiser neural network $D_{\theta}$ is trained by
@@
After obtaining a well-trained conditional diffusion model, we leverage it for sampling data and store data in augmentation dataset $\mathcal{D}^{*}$ for further policy training.
IV-F Hyperparameters
Figure 3: An illustrative example of how GODA utilizes higher goals to steer the sampling process toward higher-reward data distribution region.
TABLE I: Hyperparameter settings for denoiser network.
Hyperparameter Value
embedding dimension 128
MLP width 512
MLP activation SiLU
gate activation Sigmoid
learning rate 0.0003
batch size 256
learning rate schedule cosine annealing
optimizer Adam
gradient update steps 100K
TABLE II: Hyperparameter settings for the diffusion model.
Hyperparameter Value
number of diffusion steps 128
$S_{\text{churn}}$ 80
$S_{\text{tmin}}$ 0.05
$S_{\text{tmax}}$ 50
$S_{\text{noise}}$ 1.003
$\sigma_{\text{min}}$ 0.002
$\sigma_{\text{max}}$ 80
IV-F1 Denoising Network
The denoising neural network utilizes the adaptive gated conditioning architecture, as shown in Figure2. Table I details the associated hyperparameters. We use Random Fourier Feature embedding [30] and Sinusoidal positional embedding [31] to process the noise level and timestep of each RTG respectively, with an embedding dimension of 128. The width of the linear layers in the MLP block is set to 512, with SiLU [35] as the activation function. The total number of trainable parameters for the denoiser neural network is approximately 3.3M. We train our GODA model with 100K steps of gradient updates, with a batch size of 256 and a learning rate of 0.0003.
IV-F2 Elucidated Diffusion Model
We adopt EDM [15] as our diffusion model and follow the original settings from SynthER [13], with the default hyperparameters shown in Table II. EDM employs Heun’s $2^{\text{nd}}$ order ODE solver [36] to solve the reverse-time ODE, enabling data sampling through the reverse process. The diffusion timestep is set to 128 for higher-quality results. All training and sampling are conducted on an AMD Ryzen 7 7700X 8-Core Processor and a single NVIDIA GeForce RTX 4080 GPU. Training GODA for 100K steps takes approximately 14 minutes while generating 5M samples with a sampling batch size of 250K requires about 300 seconds.
V Experimental Settings
@@
We adopt three popular Mujoco locomotion tasks from Gym  [^1] https://www.gymlibrary.dev/environments/mujoco/, i.e., HalfCheetah, Hopper, and Walker2D, and a navigation task, i.e., Maze2D [37], as well as more complex tasks, specifically the Pen and Door tasks from the Adroit benchmark [38, 37].
V-A1 Gym
In the case of Gym tasks where dense rewards are available, we employ three distinct data configurations from the D4RL datasets: Random, Medium-Replay, and Medium. To elaborate, Random datasets contain transition data generated by a randomly initialized policy. Medium datasets consist of a million data points gathered using a policy that achieves one-third of the performance of an expert approach. Medium-replay datasets contain the stored experience in a replay buffer during the training of a policy until it reaches the score in Medium datasets.
V-A2 Maze2D
For Maze2D, a 2D agent is trained to reach a goal position utilizing minimal feedback i.e., a single point for success, zero otherwise. Three datasets collected from different maze layouts are adopted, i.e., Umaze, Medium, and Large.
V-A3 Adroit
Besides, we further test our GODA on more complex tasks, specifically the Pen and Door tasks from the Adroit benchmark [38, 37]. These tasks involve manipulating a pen and opening a door using a 24-DoF simulated hand robot. For each task, we use two different datasets: Human and Cloned. The Human dataset consists of trajectories from human demonstrations, while the Cloned dataset is generated by applying an imitation policy trained from a mix of human and expert demonstrations and combining the resulting trajectories with the demonstrations in a 50/50 split.
Figure 4: A standard signalized intersection with four three-lane approaches and eight phases.
@@ V-A3 Adroit
Figure 5: Data quality evaluation for SynthER and GODA on Walker2D-Random-V2. Left: Dymanics MSE and L2 Distance comparison. Smaller Dynamics MSE indicates better validity and larger L2 Distance indicates higher diversity. Middle: Ground-truth reward distributions from the simulator for augmented datasets. Right: Ground-truth and augmented reward distributions for GODA dataset.
@@
TABLE III: Data quality evaluation metrics for SynthER and GODA on Walker2D tasks. Smaller Dynamics MSE, larger L2 Distance, and larger Average Reward indicate better quality.
Task Dynamics MSE L2 Distance Average Reward
@@
Walker2D-Random-v2 2.7±5.7 1.9±2.9 21.8±7.0 23.2±7.6 0.1±0.6 0.6±0.5
Walker2D-Medium-Replay-v2 0.5±1.7 0.4±1.1 17.3±6.3 17.2±6.0 2.5±1.3 3.5±0.9
Walker2D-Medium-v2 0.3±1.0 0.3±0.8 11.7±5.2 11.8±5.3 3.4±1.2 3.7±0.9
V-B Traffic Signal Control Tasks and Datasets.
To evaluate GODA’s applicability to real-world challenges, we further test it on TSC tasks with much fewer training samples using the CityFlow simulator [39]. TSC aims to optimize traffic flow by efficiently managing traffic signals to maximize overall traffic efficiency. As shown in Figure 4, a signalized intersection in TSC problems is composed of approaches with several lanes in each approach. The controller manages the phase as shown in the top right part of Figure 4, which determines the activated traffic signals for different directions, to control the orderly movement of vehicles.
To evaluate our GODA, we select three real-world scenarios featuring a 12-intersection grid from Jinan (JN) city and two scenarios with a 16-intersection grid from Hangzhou (HZ) city [6]. These scenarios represent a variety of traffic patterns and intersection structures, allowing us to cover a wide range of traffic situations. To bridge the gap between simulation and real-world conditions, we use the widely adopted Fixed-Time (FT) controller as one of our behavior policies for generating the initial offline datasets. Additionally, we employ Advanced Max Pressure (AMP) [40] and Advanced CoLight (ACL) [40] to create higher-quality datasets for further evaluation.
V-C Baseline Methods
TATU [^2] https://github.com/pipixiaqishi1/TATU [11], which learns world models to generate synthetic rollouts and truncates trajectories with high accumulated uncertainty.
SynthER [^3] https://github.com/conglu1997/SynthER [13], which employs diffusion models to unconditionally augment large amounts of new data based on the learned distribution of original datasets.
DiffStitch [^4] https://github.com/guangheli12/DiffStitch [4], which augments data with a diffusion model and three MLPs, and actively connects low to high-reward trajectories with stitch techniques.
V-D Evaluation Algorithms
To verify the quality of datasets augmented by GODA, we follow the evaluation settings adopted in previous data augmentation studies. We train two widely-used offline RL algorithms, i.e., IQL [10] and TD3+BC [41], on datasets and evaluate the learned policy on D4RL tasks. For TSC tasks, we utilize BCQ [42], CQL [21], and DataLight [6] as the evaluation algorithms.
It is important to note that for GODA, we train the evaluation algorithms using a mix of the original datasets and the augmented datasets, whereas for the other baseline methods, only the augmented datasets are used. This is because GODA focuses on augmenting samples from the high-reward zones, which may lead to reduced data diversity. In contrast, the baseline methods, as reported in their respective papers [13] and our experiments, exhibit degraded or similar performance when using a mix of the original and augmented datasets, as illustrated in Section VII-C5.
V-E Implementation Details
V-E1 D4RL Tasks
We augment 5M samples for each D4RL task. For the sampling process, we use the return-prior goal condition selection method and set the goal scaling factor to 1.1 for all tasks. We use the implementation of IQL and TD3+BC from the Clean Offline Reinforcement Learning (CORL) codebase [43] for D4RL tasks.
V-E2 TSC Tasks
We formulate the TSC problem as a MDP and define the state, action, and reward function as follows:
State. For behavior policies (AMP and ACL), the state representation includes the current phase, traffic movement efficiency pressure, and the number of effective running vehicles [40]. For evaluation algorithms, BCQ and CQL use the same state representation as AMP, while DataLight adopts the number of vehicles, along with the total velocity saturation and unsaturation degrees [6].
Reward. AMP, BCQ, CQL and DataLight use pressure [40] as the reward while ACL uses queue length. It is worth noting that we use the opposite of these metrics as the final reward function.
Due to the absence of certain key signals in the original datasets from the DataLight codebase, we generate a total of 24K samples for each dataset using the behavior policies for each task. Additionally, we augment 20K samples for each task using our GODA model. The task horizon for each TSC scenario is set to 3600 seconds, with a control step of 15 seconds. For the sampling process, we employ the return-prior goal condition selection method and set the goal scaling factor to 1.2 in TSC tasks. For evaluation methods, we employ the implementation of BCQ, CQL, and DataLight from https://github.com/LiangZhang1996/DataLight.
VI Augmented Data Quality Measurement

### v2

@@ 3.2 Selective goal conditions
3.3 Controllable goal scaling
Selective goal conditions offer high-return guidance during the sampling process, but are limited in generating data with returns or quality beyond the initial offline datasets. To overcome this limitation, we introduce a controllable goal scaling factor, $\lambda$, which can be multiplied with the goal values to represent a higher return expectation. This approach enables flexible adjustment of goal values to drive the sampling process toward higher-quality data. As illustrated in Figure 1, a higher RTG goal at each timestep directs the sampling process toward a trajectory with a greater overall return. Since RTG values can be either positive or negative in certain tasks, we propose multiplying positive goals by the scaling factor and dividing negative goals by it.
It is worth noting that we use a goal scaling factor rather than an increment factor, i.e., $\text{goal}=(\hat{g}_{t}+\lambda,t)$. The reason is that incrementing each RTG $\hat{g}_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$ with a fixed value would be the same as simply incrementing the terminal reward $r_{T}$ by $\lambda$ while leaving all other target rewards $r_{t}$ for $t\neq T$ unchanged. However, assigning a higher RTG goal at each timestep is intended to encourage the agent to generate trajectories that yield higher rewards throughout the trajectory, rather than only improving the terminal reward. Therefore, a goal scaling factor can be a better choice than an increment factor.
3.4 Adaptive Gated Conditioning
3.4.1 Condition Embedding
The noise level $\sigma$ for diffusion is encoded using Random Fourier Feature (Rahimi & Recht, 2007) embedding. The RTG is processed with a linear transformation to get a hidden representation. The timestep of each RTG is embedded with Sinusoidal positional embedding (Vaswani et al., 2017). We concatenate the RTG and timestep embeddings to form the goal condition, which is then element-wise added to the noise level embedding and used as the conditional input.
3.4.2 Adaptive Gated Long Skip Connection
As shown in the left part of Figure 3, we adopt a long skip connection similar to U-Net (Ronneberger et al., 2015) to connect MLP blocks at different levels. To capture different information with varying importance weights, we propose a novel adaptive gated long skip connection structure by adding the previous information with an adaptive gate mechanism based on the given conditions.
$x_{\text{out}}=(1-\omega)*x_{\text{skip}}+\omega*x,$ (7)
where $x_{\text{skip}}$ and $x$ are outputs of a shallower block and the previous block, and $\omega$ denotes a learnable weight calculated by regressing the conditional input with an MLP and a sigmoid layer.
Figure 3: The denoiser neural network with adaptive gated conditioning architecture. The condition-adaptive gated long skip connection, shown in the left part, captures both low- and high-level features, assigning varying importance weights to each. The orange arrows to the left of each Gated Residual MLP Block represent the noise level and goal condition. The adaptive gated residual, depicted in the right part, further enhances the model by selectively preserving input information based on the given conditions.
3.4.3 Gated Residual MLP Block
The right part of Figure 3 depicts the structure of each gated residual MLP block. We mainly use a simple MLP structure to process the inputs, given the non-sequence input format. We have also tried more complicated structures, e.g., self-attention and cross-attention mechanisms, while getting degraded performance. We adopt the widely used adaptive layer normalization (adaLN) method (Peebles & Xie, 2023) to learn dimension-wise scale $\gamma$ and shift $\beta$ based on the conditional information. Besides, we explore a modification of the residual connection (He et al., 2016) and introduce a novel condition-adaptive gated residual connection to further enhance the model by selectively preserving input information. It also regresses the conditional input and gets a learnable weight $\nu$ for adaptively preserving input information.
$x_{\text{out}}=(1-\nu)*F(x)+\nu*x,$ (8)
where $F$ is the learned transformation.
@@ 3.4.3 Gated Residual MLP Block
3:Split initial offline dataset $\mathcal{D}$ into trajectories according to episode terminal information
4:Calculate RTGs for each transition sample in trajectories
5:Add RTGs and timesteps as goals into $\mathcal{D}$
6:Train $G_{\theta}$ on $\mathcal{D}$ using Eq.  9 by conditioning on goals
7:# Data Augmentation (Sampling)
8:repeat
9:  Extract goal conditions for sampling according to the goal selection mechanism.
@@
10:  Re-assign sampling goal conditions with goal scaling factor $\lambda$ using Eq. 6
@@
11:  Sampling a batch of new transition samples $\mathcal{B}^{*}$ by conditioning on scaled goals
12:  $\mathcal{D}^{*}\leftarrow\mathcal{D}^{*}\cup\mathcal{B}^{*}$
13:until $|\mathcal{D}^{*}|=$ number of target samples
14:# Policy Training and Evaluation
15:Train policy $\pi$ on the final dataset $\mathcal{D}^{*}\cup\mathcal{D}$
16:Evaluate policy on target tasks
3.5 Method Implementation
@@
Given the strong ability of diffusion models to capture the complex data distribution and generate high-dimensional data, we adopt EDM (Karras et al., 2022) as our generative model for augmenting offline data. The neural network equipped with adaptive gated conditioning, as illustrated in Figure 3, is used as the denoiser function. It takes as input the noised sample $\mathbf{x}+\mathbf{n}$, noise level $\sigma$, and goal condition $\mathbf{c}$, and make predictions for the original sample $\mathbf{x}$. Therefore, with goal conditions $\mathbf{c}$ and transition data $\mathbf{x}$ from the original datasets, the diffusion model $G_{\theta}$ with a learnable denoiser neural network $D_{\theta}$ is trained by
@@
Algorithm 1 outlines the overall process of our GODA method. We begin by training a conditional diffusion model to approximate the data distribution of the offline dataset (Lines 2-5), using transition tuples paired with their corresponding goal conditions as training samples. Once the diffusion model is well trained, we use it to generate new transition samples by conditioning it on selected and scaled goals (Lines 6-11). The resulting augmented samples are then stored in the augmentation dataset $\mathcal{D}^{*}$ for subsequent policy training (Line 12). The learned policy is finally evaluated on target tasks (Line 13).
4 Experimental Settsings
@@ 4.1.1 D4RL Benchmark
We adopt three popular Mujoco locomotion tasks from Gym  [^1] https://www.gymlibrary.dev/environments/mujoco/, i.e., HalfCheetah, Hopper, and Walker2D, and a navigation task, i.e., Maze2D (Fu et al., 2020), as well as more complex tasks, specifically the Pen and Door tasks from the Adroit benchmark (Rajeswaran et al., 2017; Fu et al., 2020). For locomotion tasks, we adopt four data quality levels: Random, Medium-Replay, Medium, and Medium-Expert. For Maze2D, three datasets collected from different maze layouts are adopted, i.e., Umaze, Medium, and Large. For the Adroit benchmark, we use two different datasets: Human and Cloned.
4.1.2 V-D4RL Benchmark
We further extend our approach to the Cheetah-Run task from V-D4RL Lu et al. (2023), a high-dimensional pixel-based offline RL benchmark, with four data quality levels: Random, Medium-Replay, Medium, and Medium-Expert. We implement our GODA by following the same settings adopted in SynthER Lu et al. (2024). Specifically, the original image-based datasets are first used to pre-train a Behavior Cloning (BC) policy network composed of a CNN encoder, a trunk network, and a final fully connected layer. The CNN encoder and trunk network are then frozen and used to transform each observation and next observation, represented as raw image data of size 84×84×3, into a 50-dimensional latent representation. These low-dimensional representation data, combined with rewards and actions, are then used to train the diffusion model and augment transition samples in the latent space. These augmented samples are subsequently used to fine-tune only the final fully connected layer (output head), thereby avoiding the need to directly generate high-dimensional image data.
4.1.3 Traffic Signal Control
To evaluate GODA’s applicability to real-world challenges, we further test it on TSC tasks with much fewer training samples using the CityFlow simulator (Zhang et al., 2019). TSC aims to optimize traffic flow by efficiently managing traffic signals to maximize overall traffic efficiency.
To evaluate our GODA, we select three real-world scenarios featuring a 12-intersection grid from Jinan (JN) city and two scenarios with a 16-intersection grid from Hangzhou (HZ) city (Zhang & Deng, 2023). These scenarios represent a variety of traffic patterns and intersection structures, allowing us to cover a wide range of traffic situations. To bridge the gap between simulation and real-world conditions, we use the widely adopted Fixed-Time (FT) controller as one of our behavior policies for generating the initial offline datasets. Additionally, we employ Advanced Max Pressure (AMP) (Zhang et al., 2022) and Advanced CoLight (ACL) (Zhang et al., 2022) to create higher-quality datasets for further evaluation. We present more details in Appendix B.1.4.
4.2 Baseline Data Augmentation Methods
TATU [^2] https://github.com/pipixiaqishi1/TATU (Zhang et al., 2023), which learns world models to generate synthetic rollouts and truncates trajectories with high accumulated uncertainty.
SynthER [^3] https://github.com/conglu1997/SynthER (Lu et al., 2024), which employs diffusion models to unconditionally augment large amounts of new data based on the learned distribution of original datasets.
DiffStitch [^4] https://github.com/guangheli12/DiffStitch (Li et al., 2024), which augments data with a diffusion model and three MLPs, and actively connects low to high-reward trajectories with stitch techniques.
4.3 Evaluation Algorithms
To verify the quality of datasets augmented by GODA, we follow the evaluation settings adopted in previous data augmentation studies. We train two widely-used offline RL algorithms, i.e., IQL (Kostrikov et al., 2021) and TD3+BC (Fujimoto & Gu, 2021), on datasets and evaluate the learned policy on D4RL tasks. For TSC tasks, we utilize BCQ (Fujimoto et al., 2019), CQL (Kumar et al., 2020), and DataLight (Zhang & Deng, 2023) as the evaluation algorithms. For V-D4RL tasks, the BC algorithm is adopted for evaluation.
((a))
((b))
((c))
@@ 4.3 Evaluation Algorithms
Figure 4: Data quality evaluation for SynthER and GODA on Walker2D-Random-V2. Left: Dynamics MSE and Min L2 Distance comparison. Smaller Dynamics MSE indicates better validity, and larger Min L2 Distance indicates higher diversity. Middle: Ground-truth reward distributions from the simulator for augmented datasets. Right: Ground-truth and augmented reward distributions for the GODA-augmented dataset.
It is important to note that for GODA, we train the evaluation algorithms using a mix of the original datasets and the augmented datasets, whereas for the other baseline methods, only the augmented datasets are used. This is because GODA focuses on augmenting samples from the high-reward zones, which may lead to reduced data diversity. In contrast, the baseline methods, as reported in their respective papers (Lu et al., 2024) and our experiments, exhibit degraded or similar performance when using a mix of the original and augmented datasets, as illustrated in Section C.5.
@@
Table 1: Data quality evaluation metrics for SynthER and GODA on Walker2D tasks. Smaller Dynamics MSE, larger Min L2 Distance, and larger Average Reward indicate better quality.
Task Dynamics MSE Min L2 Distance Average Reward
@@
Walker2D-Random-v2 2.7±5.7 1.9±2.9 1.9±0.6 2.4±0.6 0.1±0.6 0.6±0.5
Walker2D-Medium-Replay-v2 0.5±1.7 0.4±1.1 1.6±0.9 1.7±0.7 2.5±1.3 3.5±0.9
Walker2D-Medium-v2 0.3±1.0 0.3±0.8 0.8±0.5 0.8±0.6 3.4±1.2 3.7±0.9
5 Augmented Data Quality Measurement
Since our GODA is built upon SynthER, we compare the quality of the datasets augmented by both SynthER and GODA to assess whether the goal conditions incorporated by GODA enhance data quality. We adopt two metrics, i.e., Dynamics MSE and Min L2 Distance, from SynthER (Lu et al., 2024):
$\text{Dynamics MSE}=\frac{1}{M}\sum_{i=1}^{M}\left((s_{t+1}^{i},r_{t}^{i})-(\hat{s}_{t+1}^{i},\hat{r}_{t}^{i})\right)^{2},$ (10)
[…]
$\text{Average Reward}=\frac{1}{M}\sum_{i=1}^{M}\hat{r}_{t}^{i},$ (12)
Dynamics MSE measures how well the augmentation models capture the dynamics of the environment by learning patterns from the original datasets and generating data that aligns with those dynamics. Min L2 Distance assesses the models’ exploration capabilities and data diversity by calculating the average L2 distance between each augmented sample and its nearest neighbor in the original dataset, reflecting how diverse the generated data is. Average Reward compares the ground-truth reward distributions produced by the simulator given states and actions in datasets augmented by SynthER and GODA.
The left part of Figure 4 presents a scatter plot of 10K points sampled from the augmented datasets. Results show that datasets generated by GODA exhibit much lower Dynamics MSE and a wider range of Min L2 Distance values, indicating both better alignment with environmental dynamics and greater diversity. The superior validity in performance may stem from the goal conditions (RTG-timestep pairs), which provide critical information for generating samples that better match the environment’s dynamics. Meanwhile, the increased diversity is likely due to the scaled-out-of-distribution goal conditions incorporated in the sampling process.
The evaluation results for the three metrics in Table 1 further demonstrate that GODA outperforms SynthER in terms of all data quality evaluation metrics across nearly all Walker2D tasks. Notably, GODA is particularly effective on random-level datasets with high randomness and sparse optimal demonstrations, where traditional augmentation methods often fall short.